OpenStudy (anonymous):

is the common denominator for 4/5+1/4 20?

jimthompson5910 (jim_thompson5910):

Yes, this is because the LCM of 5 and 4 is 20.

OpenStudy (anonymous):

4*4=16 5*4=20 16/20+5/20=21/20?